OpenStudy (anonymous):

Solve Z = 9x + 9y for y.

OpenStudy (anonymous):

Solving for y means that we want y to be on one side by itself so let's keep it on the left side Z = 9x + 9y Let's get rid of that 9x then. To do that we need to subtract 9x to both sides. What do we get then?

OpenStudy (wach):

First, isolate y to one side. You can do this by subtracting 9x from both sides, and then further ridding 9y of 9 by dividing by 9.

OpenStudy (anonymous):

Sorry I meant right side.

OpenStudy (anonymous):

Z= 9y

OpenStudy (anonymous):

wait didn't you subtract 9x on the left side?

OpenStudy (anonymous):

-9x + z = 9y

OpenStudy (anonymous):

great now we divide both sides by 9 . This will get rid of that 9 next to the y

OpenStudy (anonymous):

-9x + z/9 =9y/9

OpenStudy (anonymous):

y= -9x + z /9

OpenStudy (anonymous):

yes\[\frac{-9x + z}{9} =y\]

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

could it also be written as y = z - 9x/9?

OpenStudy (anonymous):

Thanks for the help with problem